Paladin Furnishings generated $2 million in sales during 2016, and its year-end total assets were $1.7 million. Also, at year-end 2016, current liabilities were $500,000, consisting of $200,000 of notes payable, $200,000 of accounts payable, and $100,000 of accrued liabilities. Looking ahead to 2017, the company estimates that its assets must increase by $0.85 for every $1.00 increase in sales. Paladin's profit margin is 5%, and its retention ratio is 40%.

How large of a sales increase can the company achieve without having to raise funds externally?
Answer #1

Sales in 2016, S2016 = $ 2,000,000

Assets in 2016, A2016 = $ 1,700,000

Current Liabilities in 2016, CL2016 = $ 500,000

Notes payable in 2016, NP2016 = $ 200,000

Accounts payable in 2016, A/P2016 = $ 200,000

Accruals in 2016, A2016 = $ 100,000

Profit Margin, M = 5 %

Retention ratio, r = 40 %

Increase in assets w.r.t. sales, A*/S0 = 0.85

Additional fund needed, AFN = (A*/S0) ΔS – (L*/S0) ΔS – M x S1 x r

                                             = 0.85 x ΔS – [($ 200,000 + $ 100,000/S0)] ΔS – 0.05 x S1 x 0.4

                                             = 0.85 x ΔS – ($ 300,000/$ 2,000,000) ΔS – 0.05 x S1 x 0.4

                                            = 0.85 x ΔS – 0.15 x ΔS – 0.02 x S1

                                                           = 0.70 x ΔS – 0.02 x S1

                                            = 0.70 x (S1 – S0) – 0.02 x S1

                                            = 0.70 x (S1 – $ 2,000,000) – 0.02 x S1

                                            = 0.7 x S1 – 0.7 x $ 2,000,000 – 0.02 x S1

                                            = 0.7 x S1 – $ 1,400,000 - 0.02 x S1

                                            = 0.68 S1 – $ 1,400,000

                       IF AFN = 0;

                                 0.68 S1 = $ 1,400,000

                                S1 = $ 1,400,000/0.68 = $ 2,058,823.53

Increase in sales = S1 – S0 = $ 2,058,823.53 - $ 2,000,000 = $ 58,823.53

Company can achieve a sales increase by $ 58,823.53 without raising fund externally.